# Predictable Scale: Part I — Optimal Hyperparameter Scaling Law in Large Language Model Pretraining

### v4

@@ Abstract
The impressive capabilities of Large Language Models (LLMs) across diverse tasks are now well-established, yet their effective deployment necessitates careful hyperparameter optimization. Through extensive empirical studies involving grid searches across diverse configurations, we discover universal scaling laws governing these hyperparameters: optimal learning rate follows a power-law relationship with both model parameters and data sizes, while optimal batch size scales primarily with data sizes. Our analysis reveals a convex optimization landscape for hyperparameters under fixed models and data size conditions. This convexity implies an optimal hyperparameter plateau. We contribute a universal, plug-and-play optimal hyperparameter tool for the community. Its estimated values on the test set are merely 0.09% away from the globally optimal LLM performance found via an exhaustive search. These laws demonstrate remarkable robustness across variations in model sparsity, training data distribution, and model shape. To our best known, this is the first work that unifies different model shapes and structures, such as Mixture-of-Experts models and dense transformers, as well as establishes optimal hyperparameter scaling laws across diverse data distributions. This exhaustive optimization process demands substantial computational resources, utilizing nearly one million NVIDIA H800 GPU hours to train 3,700 LLMs of varying sizes and hyperparameters from scratch and consuming approximately 100 trillion tokens in total. To facilitate reproducibility and further research, we will progressively release all loss measurements and model checkpoints through our designated repository https://step-law.github.io/.
@@ 1 Introduction
Name Data Recipe Model Sparsity LR BS Relative Error
OpenAI Law
[…]
Microsoft Law
[…]
DeepSeek Law
[…]
Porian Law
[…]
MiniCPM Law
@@
Table 1: Comparison of optimal hyperparameter scaling laws across different approaches. Data Recipe and Model Sparsity denotes whether the approach is suitable for different data recipe and model sparsity. Relative Error denotes the relative loss, as same as Fig 1. The variables in scaling laws are described in Section 1.1.
State-of-the-art Large Language Models (LLMs) Brown et al. (2020); Jin et al. (2023); Biderman et al. (2023); Scao et al. (2022); Touvron et al. (2023a, b); Grattafiori et al. (2024); DeepSeek-AI et al. (2024a); Yang et al. (2024); DeepSeek-AI et al. (2024b, 2025), have reached unprecedented scales, with models being trained on billions of parameters and trillions of tokens. Recent developments like Llama 3 Grattafiori et al. (2024) demonstrate this trend, utilizing 15 trillion tokens for training Grattafiori et al. (2024). At such massive scales, identifying optimal hyperparameter configurations becomes both critical for model performance and challenging due to computational constraints.
Figure 1: This plot shows the hyperparameter space for a model with 1 billion (B) parameters trained on 100B tokens. We trained 120 LLMs from scratch with different learning rate (LR) and batch size (BS) combinations, obtaining contour lines and global optimal points based on real data. Optimal points represent the lowest training loss for each LR and BS pair, while contour lines depict the relative loss differences from these optima. Our Step Law predicts the optimum with the highest accuracy compared to other methods, nearly matching the global optimal points.
The success of LLM pretraining heavily depends on hyperparameter settings, particularly the learning rate and batch size. Suboptimal configurations can lead to various issues: excessive learning rates may cause training divergence, while insufficient rates slow down progress Shen et al. (2024); Wen et al. (2024); similarly, batch size must balance computational efficiency and model quality Perko (2023); Filatov et al. (2024); McCandlish et al. (2018). Traditional grid search becomes prohibitively expensive at scale, leading researchers to rely on hyperparameter transfer methods extrapolating optimal configurations from smaller-scale experiments to larger ones Yang and Hu (2020); Yang et al. (2023).
@@
Prior work in hyperparameter transfer broadly falls into two categories: theory-driven and data-driven approaches. In theory-driven methods, $\mu$P Yang et al. (2022) pioneered learning rate transfer rules across model widths, with subsequent work Everett et al. (2024); Lingle (2024); Blake et al. (2024); Yang et al. (2023); Bordelon et al. (2023) extending these findings to various model depths while also revealing their limitations. In data-driven approaches, Kaplan et al. (2020) established foundational learning rate scaling laws based on model size $N$, inspiring further investigations Bjorck et al. (2024); DeepSeek-AI et al. (2024a); Porian et al. (2024); Hu et al. (2024) into learning rate (LR) and batch size (BS) scaling for dense models. Recently, Wang et al. (2024); Ludziejewski et al. (2025) had begun exploring these relationships in Mixture-of-Experts (MoE) modelsDu et al. (2021); Fedus et al. (2021).
However, a significant gap remains in understanding hyperparameter transfer across different dimensions: data recipe, model shape, model sparsity, and data sizes $D$. While existing research has made progress in understanding scaling behavior across model sizesKaplan et al. (2020); Halfon et al. (2024) the interaction of these other critical factors remains under-explored. Our work addresses this gap by empirically discovering universal hyperparameter (HP) scaling laws that hold across these varied dimensions, providing a more comprehensive understanding of optimal hyperparameter selection in LLM pretraining.
@@
Our main contributions are as follows:
(i) We are the first to discover and demonstrate the convexity property of the loss landscape under fixed parameter count and data size conditions. This provides fundamental insights into hyperparameter optimization, as shown in Fig. 2.
(ii) We establish the first universal and robust hyperparameter scaling laws for LLM pretraining, which is called Step Law. Our Empirically discovered the power-law relationship between optimal learning rate $\eta(N,D)$ and optimal batch size $B(D)$. Step Law demonstrates that the optimal batch size exhibits a primary dependence on dataset size $D$, while the optimal learning rate manifests a joint dependence on both model parameters $N$ and dataset size $D$. Step Law is defined as Eq.(1).
The Step Law achieves substantially superior convergence results compared to baseline methods when generalized to 1 billion parameter models, as illustrated in Fig. 1. Step Law provides a plug-and-play formula that eliminates extensive hyperparameter tuning efforts for industry applications.
(iii) We are the first to study the transferability and invariance of optimal hyperparameter scaling laws across different pretraining data distributions. We systematically analyze how optimal hyperparameter scaling laws transfer across different pretraining data distributions and model architectures. Our work pioneers the investigation into whether dense LLMs and sparse (MoE) LLMs with varying sparsity levels share common optimal hyperparameter patterns, revealing significant invariance properties between them. Through extensive grid search experiments, we validate that Step Law maintains high generalizability and robustness across different pretraining corpora distributions, model architectures, and both dense and sparse (MoE) LLMs with varying sparsity ratios.
(iv) We conduct an unprecedented large-scale empirical study, involving:
  • Extensive experimentation across 3700 model configurations, training LLMs from scratch with Dense and MoE (varying sparsity ratios), model architectures, data distributions, and hyperparameter settings.
  • Total compute consumption approaching 1 million H800 GPU Hours (equivalent to over \$1 million), processing approximately 100 trillion tokens during training.
This represents the largest dataset of hyperparameter optimization results in the field, derived purely from empirical observations without prior assumptions. Training checkpoints and hyperparameter configurations will be made publicly available.
1.1 Notation
We use the following notation:
  • $\mathcal{L}$: Cross-entropy loss.
  • $D$: Dataset size in tokens.
  • $N$: Number of non-embedding parameters in the model.
@@ 1.1 Notation
  • $C$: Compute budget in FLOPs.
  • $N_{{layer}}$: Number of layers in the Transformer model.
  • $d_{ff}$: Dimension of the feed-forward network hidden layer in the Transformer.
  • $d_{{model}}$: Hidden dimension of the Transformer model.
  • $N_{{head}}$: Number of attention heads in the Transformer model.
  • $B(N,D)$: Optimal batch size (in tokens) for a given parameter count $N$ and dataset size $D$.
2 Related Works
Hyperparameter transfer, which involves extrapolating optimal settings from smaller to larger models, has become essential for efficient large-scale training. Among these, learning rate (LR) and batch size (BS) are particularly crucial hyperparameters that substantially influence LLM pretraining performanceHalfon et al. (2024). Research on optimal learning rate and batch size selection broadly falls into two categories: theory-driven and data-driven approaches.
In theory-driven approaches, $\mu\text{P}$ Yang et al. (2022) established foundational learning rate transfer rules for varying model widths, though this required specific modifications to initialization and attention mechanisms. However, the $\mu\text{P}$ framework and its extensions Yang et al. (2022); Everett et al. (2024); Lingle (2024); Blake et al. (2024); Yang et al. (2023); Bordelon et al. (2023) are limited in scope, lacking guidance for learning rate adaptation across different data distributions, model architectures, sparsity levels, and data sizes. Additionally, these works do not address optimal batch size prediction.
In data-driven approaches, a fundamental principle in deep learning is that larger models require smaller learning rates to ensure training stability and convergence. Kaplan et al. (2020) formalized this relationship, expressing learning rate as a function of model size. Bjorck et al. (2024) incorporated data size dependency by proposing $LR(N,D)=CN^{-\alpha}D^{-\beta}$. Batch size optimization is equally important for balancing convergence and computational efficiency. While several approaches followed Kaplan et al. (2020)’s framework, they face limitations: Wang et al. (2024); Hu et al. (2024) derived batch size based on expected loss but require prior knowledge of model behavior; Porian et al. (2024) refined scaling laws across two datasets but only considered model size, setting final learning rates at 0.1% of peak values.
Notably, both early work McCandlish et al. (2018) and recent Critical Batch Size (CBS) analysis Zhang et al. (2024) support our empirical finding that optimal batch size primarily depends on dataset size rather than model size, with CBS establishing a theoretical upper bound for this relationship. DeepSeek-LLM DeepSeek-AI et al. (2024a), while comprehensive in its approach using IsoFLOP Hoffmann et al. (2022), is constrained by its fixed compute budget assumption.
Our research advances these findings by demonstrating that model size and dataset size are sufficient predictors of optimal hyperparameters. We validate these scaling rules across diverse architectures, including variations in model sparsity, data distributions and model shape.
Figure 3: Contour plots of training loss (left) and validation loss (right) across hyperparameter configurations. Both plots share the global minimum () at batch size 393,216 and the learning rate of 0.001950.
3 Problem Setup
3.1 Preliminaries
@@ 3.2 Experimental Settings
We train our models using language modeling loss on a dataset that includes web text, mathematical content, and code. The dataset is tokenized using a BPE Gage (1994) tokenizer with a vocabulary size of 65,536.
[…]
@@
We use the AdamW Loshchilov and Hutter (2017) optimizer with $\beta$ values of [0.9, 0.95], an epsilon of $10^{-8}$, a weight decay of 0.1, and a gradient clipping norm of 1.0. Our learning rate schedule includes a linear warmup phase over the initial 2,000 steps, followed by a cosine decay reaching a final learning rate of $10^{-5}$ for the remainder of the training. A detailed analysis and rationale for this strategy are provided in Section 4.1.3. The sequence length is set to 2,048 tokens. We conduct experiments using training data proportions aligned with Llama-2  Touvron et al. (2023b) ( More details are described in Tab. 6 ).
The learning rate is selected from a logarithmic sequence of powers of 2, spanning exponents from -10.5 to -7.0 in regular 0.5-interval increments. The batch size is selected from a predefined geometric progression ranging from 32,768 to 4,194,304, where each subsequent batch size is obtained by multiplying the previous value by a constant factor of 2, maintaining an exponential growth trend. Both parameter configurations correspond to the 18 LLMs detailed in Tab. 4 in Appendix A.2.
4 Experiments
4.1 Ablations
4.1.1 Evaluation metric
As described in Chinchilla Hoffmann et al. (2022), smoothed training loss is considered an unbiased estimate of validation loss for simplicity. We operate under this same setting and supplement our investigation with experimental analysis.
As shown in Fig. 3, for the case where the smoothed training loss converges to the optimal value of $2.279$ ( as indicated by the solid red-framed line in Fig. 3 right) , the corresponding LR and BS are $1.95\times 10^{-3}$ and $393,216$ respectively. This is the same as the position of the LR and BS corresponding to the validation loss converging to the optimal value of $2.038$ (as indicated by the solid red-framed line in Fig. 3 left). Moreover, the overall trend of how the smoothed training loss deviates from the optimal value with varying learning rates and batch sizes (as shown by the heatmap patterns in Fig. 3 right) closely mirrors the corresponding variations observed in validation loss measurements. This alignment demonstrates that the smoothed training loss provides consistent optimization guidance for learning rate and batch size selection, matching the parameter configurations that would be obtained through direct validation loss evaluation.
4.1.2 Loss Landscape Convexity with LR and BS
To investigate the property of the loss landscape with respect to learning rate and batch size, we conducted systematic experiments across a wide range of configurations under fixed model parameters $N$ and data size $D$. As shown in Fig. 2, one of our experiment settings described in Section 3.2 demonstrates this property.
Figure 4: Comparison of learning rate schedules. These contour plots illustrate two distinct learning rate schedules. Blue contours represent the conventional decay schedule, where the minimum learning rate ($\min\_lr$) is set to one-tenth of the maximum learning rate ($\frac{\max\_lr}{10}$). Red contours depict our proposed fixed final learning rate schedule, with a constant minimum learning rate of $\min\_lr=10^{-5}$. The visualization reveals that the conventional decay method leads to a discernible leftward bias in the optimal learning rate range, indicated by the shift of the lowest loss region towards lower learning rates in the blue contours compared to the red.
@@ 4.1.2 Loss Landscape Convexity with LR and BS
Through extensive empirical analysis, we discovered a fundamental property of the loss landscape with respect to hyperparameters: both the learning rate and batch size exhibit convex relationships with the training loss under fixed model parameters and data size conditions. ( As shown in Fig. 2 one of our experiment setting described in section.)
@@
Furthermore, we observe that the loss surface demonstrates a stable region around the optimal configuration, evidenced by the plateau-like behavior shown in Fig. 3. This stability provides practical tolerance for small deviations in hyperparameter selection while maintaining near-optimal performance.
These properties form the theoretical foundation for our subsequent development of scaling laws and validate their applicability across different architectural configurations.
4.1.3 Fixed Final Learning Rate Strategy
@@
We investigated two approaches for the final minimum learning rate ($\min\_lr$): the conventional decay schedule ($\min\_lr=\frac{\max\_lr}{10}$)Brown et al. (2020); Jin et al. (2023); Touvron et al. (2023a, b); Biderman et al. (2023); Scao et al. (2022); Shen et al. (2024), and our proposed fixed schedule ($\min\_lr=10^{-5}$). Using 1B model training for 80B tokens, we compared these schedules across various LR and BS.
@@
Fig. 4 presents comparative heatmaps of the final training loss. We observe that compared to using a fixed final learning rate, setting it as max_lr/10 shows distinct optimal hyperparameter points and an overall left-skewed distribution of suboptimal learning rate and batch size combinations. We analyze that this is because, for the relatively high peak learning rates, conventional schedules result in disproportionately large minimum learning rates, which adversely affects the final stages of training and prevents the loss from converging to better local optima.
As can also be seen in Fig. 1, aside from Porian Law, which converges the min_lr to a sufficiently small value, the optimal learning rates calculated by other traditional learning rate decay schedules all exhibit varying degrees of a left-skew issue.
@@
This aligns with advanced training practices which suggest that the minimum learning rate significantly impacts the loss. This phenomenon is unfavorable for fitting our scaling laws, and in practice, it is generally preferred to keep the min_lr fixed at a relatively low value. So we adopt the fixed final learning rate strategy in our subsequent experiments.
Key Takeaways Convex Loss Landscape: The loss landscape exhibits convexity with respect to both learning rate and batch size. This convexity, coupled with a stable plateau around the optimum, underpins the robustness of hyperparameter selection. Fixed Final Learning Rate Benefit: Compared to setting a small, fixed final learning rate, the traditional decay to max_lr/10 causes the discovered optimal learning rate to be biased towards lower values (left-skewed).
4.2 Fitting HP Scaling Laws
4.2.1 LR and BS with Respect to $N$ and $D$
In accordance with Definition 3.1, we experimentally derive the LR and BS by keeping other variables fixed. This section focuses on elucidating the relationships between these empirically determined hyperparameters and $N$ and $D$. For the parameter count $N$, we set up seven experiments spanning 60M, 120M, 210M, 270M, 430M, 540M, and 1B parameters. As demonstrated in Fig. 6(a), our experiments reveal a positive correlation between optimal LR and BS and the data scale $D$ for each value of $N$. Furthermore, we conducted experiments across five different data scales $D$: 2B, 4B, 8B, 20B, and 100B tokens. Notably, we specifically reserved the 1B parameter and 100B token settings as test points to validate our findings, as discussed in Section 4.2.4. As visualized in Fig. 6(b), we find that for each data scale $D$, the optimal LR increases with model size $N$. Notably, our findings indicate that optimal BS is largely independent of $N$. Based on these experimental observations, we will present and fit the Hyperparameter (HP) scaling law formulations in Section 4.2.2.
Figure 5: The Illustration of Hyperparameter Configuration Space for 210M model parameters and 100B tokens.

### v6

# Predictable Scale: Part I, Step Law – Optimal Hyperparameter Scaling Law in Large Language Model Pre-training

@@ Abstract
The impressive capabilities of Large Language Models (LLMs) across diverse tasks are now well-established, yet their effective deployment necessitates careful hyperparameter optimization. Although existing methods have explored the influence of hyperparameters on model performance, a principled and generalizable framework across model architectures and data recipes remains absent. In this study, we conduct an unprecedented empirical investigation- training over 3,700 LLMs from scratch across 100 trillion tokens, consuming nearly one million NVIDIA H800 GPU hours to establish a universal Scaling Law for hyperparameter optimization in LLM Pre-training, called Step Law. We empirically observe that, under fixed model size ($N$) and dataset size ($D$), the hyperparameter landscape exhibits convexity with a broad optimum, substantially reducing the complexity of hyperparameter search. Building on this insight, we formally define and empirically validate the Step Law: The optimal learning rate follows a power-law relationship with $N$ and $D$, while the optimal batch size is primarily influenced by $D$ and remains largely invariant to $N$. Notably, our estimated optima deviate from the global best performance found via exhaustive search by merely 0.094% on the test set. To our best known, Step Law is the first that unifies different model shapes and structures, such as Mixture-of-Experts models and dense transformers, as well as establishes optimal hyperparameter scaling laws across diverse data recipes. We contribute a universal, plug-and-play optimal hyperparameter tool for the community, which is expected to advance efficient LLM training at scale. All experimental code, data and checkpoints are publicly available at https://github.com/step-law/steplaw.
@@ 1 Introduction
State-of-the-art Large Language Models (LLMs) have surged to unprecedented scales, training on billions of parameters and trillions of tokens Brown2020 ; Jin2023 ; biderman2023pythiasuiteanalyzinglarge ; workshop2023bloom176bparameteropenaccessmultilingual ; Touvron2023a ; Touvron2023b ; Grattafiori2024 ; DeepSeek-AI2024 ; Yang2024 ; DeepSeek-AI2024b ; DeepSeek-AI2025 . For example, Llama 3 Grattafiori2024 was trained on 15 trillion tokens, underscoring the escalating computational demands. At this scale, training hyperparameters, particularly the learning rate (LR) and batch size (BS), is extremely critical. Excessively large learning rates can cause training to diverge, whereas insufficient learning rates will result in slow convergence Shen2024 ; Wen2024 . Likewise, batch size must strike a balance between throughput and generalization Perko2023 ; Filatov2024 ; McCandlish2018 in LLM pre-training. Traditional grid search is prohibitively expensive, prompting transfer methods that extrapolate optimal settings from smaller models to large ones Yang2020 ; Yang2023 .
Name Data Recipe Model Sparsity Learning Rate Batch Size Relative Error
@@
[…]
Table 1: Comparison of optimal hyperparameter scaling laws across different approaches. Data Recipe and Model Sparsity denotes whether the approach is suitable for different data recipe and model sparsity. Relative Error denotes the relative loss, as same as Fig. 1. The variables in scaling laws are described in Sec. A.1.
To avoid exhaustive searches, the community has advanced hyperparameter-transfer rules. $\mu$P Yang2022 first established learning rate scaling with model width, later extended to depth and other architectures Everett2024 ; Lingle2024 ; Blake2024 . Empirical scaling laws then linked learning rate and batch size to model size kaplan_scaling_2020 , refined for dense models in subsequent work Bjorck2024 ; DeepSeek-AI2024 ; Porian2024 ; Hu2024 . Recently, studies have explored these interactions in Mixture-of-Experts (MoE) models Du2021 ; Fedus2021 ; Wang2024 ; Ludziejewski2025 .
Figure 1: This plot shows the hyperparameter space for a 1B model trained on 100B tokens. We trained 120 LLMs from scratch with different LR and BS combinations, obtaining contour lines and global optimal points based on real data. Optimal points represent the lowest training loss for each LR and BS pair, while contour lines depict the relative loss differences from these optima. Step Law predicts the optimum with the highest accuracy compared to other methods, nearly matching the global optimal points.
However, a significant gap remains in understanding hyperparameter transfer across different dimensions: data recipe, model shape, model sparsity, and dataset sizes $D$. Although existing research has tried to understand scaling behavior across model sizes kaplan_scaling_2020 ; Halfon2024 , the interaction of these other critical factors remains under-explored. As illustrated in Tab. 1, existing scaling laws almost uniformly fail to accommodate variations in data recipe and model sparsity. In contrast, our work is the only approach that simultaneously supports diverse data recipes and sparsity levels and reduces the relative error to merely 0.94‰, with a significant improvement over the prior methods. To address this gap, we derive universal hyperparameter scaling laws that span all key dimensions of LLM pre-training and guide optimal parameter selection. Our results show that a single empirically obtained law generalizes across model shape, sparsity, data recipe and dataset size, improving both accuracy and applicability.
Our main contributions are as follows:
(i) This paper establish the first universal and robust Scaling Law for hyperparameter optimization in LLM Pre-training, called Step Law. We discover the power-law relationship between optimal learning rate $\eta(N,D)$ and optimal batch size $B(D)$. Step Law demonstrates that the optimal batch size exhibits a primary dependence on dataset size $D$, while the optimal learning rate manifests a joint dependence on both model parameters $N$ and dataset size $D$:
@@
Step Law achieves superior convergence results compared to baseline methods when generalized to 1B models, as shown in Fig. 1. The hyperparameters predicted by Step Law yield a test-set loss within just 0.094% above the global optimum found via exhaustive search, significantly outperforming all other methods. It provides a plug-and-play formula that significantly minimizes hyperparameter tuning efforts, making it highly practical for industrial-scale training.
(ii) To the best of our knowledge, we are the first to discover and demonstrate the convexity property of the loss landscape under fixed parameter count and dataset size conditions. This significant preliminary provides fundamental insights into hyperparameter optimization, as shown in Fig. 2.
@@
(iii) Step Law is the first to investigate the transferability and robustness of optimal hyperparameter scaling laws across diverse pre-training data recipes and model architectures. We find that Step Law exhibits strong invariance across dense and sparse (MoE) LLMs with varying sparsity levels, and generalizes well across data recipes, model architectures, and sparsity configurations, as confirmed by extensive grid search. In addition, we evaluate Step Law on a 6.5B MoE model trained across varying dataset sizes. In all settings, the learning rate and batch size predicted by Step Law yield the lowest validation smooth loss, outperforming other methods.
(iv) We conduct an unprecedented large-scale empirical study, comprising 3,700 full LLM training runs across Dense and MoE models (with varying sparsity levels), diverse architectures, data recipes, and hyperparameter settings. The total compute consumed approaches 1 million H800 GPU hours (equivalent to over one million dollars), processing approximately 100 trillion tokens. To the best of our knowledge, this represents the largest dataset of hyperparameter optimization results in the field, derived entirely from empirical observations without relying on prior assumptions. All training checkpoints and hyperparameter configurations are publicly available.
2 Related Works
Hyperparameter transfer extrapolates optimal settings from smaller to larger models and is essential for efficient large-scale LLM training. In particular, learning rate and batch size have a major impact on pre-training performance Halfon2024 . Existing work divides into theory-driven and data-driven methods.
@@ 2 Related Works
Data-driven methods express hyperparameters as functions of model size $N$ and dataset size $D$. kaplan_scaling_2020 first related learning rate to $N$ , and later work incorporated $D$ into a joint law $LR(N,D)=CN^{-\alpha}D^{-\beta}$ Bjorck2024 . Batch size rules have been proposed based on expected loss Wang2024 ; Hu2024 or two-dataset fits Porian2024 , but these either require prior loss estimates, ignore $N$’s influence on learning rate, or assume fixed compute budgets DeepSeek-AI2024 . Early studies also show that optimal batch size depends mainly on $D$ rather than $N$ McCandlish2018 ; Huawei-Noah-Ark-Lab2024 ; zhang2024 .
In contrast, our Step Law demonstrates that $N$ and $D$ alone suffice to predict both learning rate and batch size. We validate these scaling laws across model shapes, sparsity levels and data recipes, offering unified and accurate hyperparameter guidance for LLM pre-training.
3 Methodology
3.1 Problem Setup
@@ 3.2 Experimental Settings
We train our models using a language modeling loss on a dataset comprising web text, math, and code, with proportions aligned with those in Llama  Touvron2023a ; Touvron2023b (see Tab. A.4 for details). The dataset is tokenized using a BPE Gage1994ANA tokenizer with a vocabulary size of 65,536.
Following the prevailing configuration in recent large-scale models, our model architecture uses RMSNorm Zhang2019 for pre-normalization and the SwiGLU Shazeer2020 activation function in the feed-forward network, without applying dropout JMLR:v15:srivastava14a . We mainly use ALiBi Press2021 positional encoding. The models are initialized from scratch, with weights sampled from a truncated normal distribution (mean of 0, standard deviation of 0.02).
@@
We use the AdamW Loshchilov2017 optimizer with $\beta$ values of [0.9, 0.95], an epsilon of $10^{-8}$, a weight decay of 0.1, and a gradient clipping norm of 1.0. Our learning rate schedule includes a linear warmup phase over the initial 2,000 steps, followed by a cosine decay reaching a final learning rate of $10^{-5}$ for the remainder of the training. A detailed analysis and rationale for this strategy are provided in Sec. 3.3.2. The sequence length is set to 2,048 tokens. The learning rate is selected from a logarithmic sequence of powers of 2, spanning exponents from -10.5 to -7.0 in increments of 0.5. The batch size is selected from a geometric progression, ranging from 32,768 to 4,194,304, with each subsequent value being $\sqrt{2}$ times the previous one. These parameter configurations correspond to the 18 LLMs detailed in Tab. 5 in Appendix A.3.
3.3 Preliminary Experiments
3.3.1 Loss Landscape Convexity Analysis
@@
Through extensive empirical analysis, we identify a fundamental property of the loss landscape with respect to hyperparameters: both the learning rate and batch size exhibit convex relationships with the training loss under fixed model parameters and dataset size conditions. Fig. 2 illustrates a representative experimental result, with comprehensive results further elaborated in Appendix A.6.
@@
Furthermore, we observe that the loss surface demonstrates a stable region around the optimal configuration, evidenced by the plateau-like behavior shown in Fig. 4. This stability provides practical tolerance for small deviations in hyperparameter selection while maintaining near-optimal performance. These properties form the theoretical foundation for our subsequent development of scaling laws and validate their applicability across different architectural configurations.
3.3.2 Impact of Final Learning Rate Schedule
Figure 3: Comparison of learning rate schedules. Contour plots illustrate two learning rate schedules: the conventional decay schedule (blue), which sets $\mathrm{lr}_{\min}=\mathrm{lr}_{\max}/10$, and our fixed schedule (red), with $\mathrm{lr}_{\min}=10^{-5}$. The visualization reveals that the conventional decay method leads to a discernible leftward bias in the optimal learning rate range, indicated by the shift of the lowest loss region towards lower learning rates in the blue contours compared to the red.
@@ 3.3.2 Impact of Final Learning Rate Schedule
We investigated two approaches for the final minimum learning rate ($\mathrm{lr}_{\min}$): the conventional decay schedule ($\mathrm{lr}_{\min}=\mathrm{lr}_{\max}/10$) Brown2020 ; Jin2023 ; Touvron2023a ; Touvron2023b ; biderman2023pythiasuiteanalyzinglarge ; workshop2023bloom176bparameteropenaccessmultilingual ; Shen2024 , and proposed fixed schedule ($\mathrm{lr}_{\min}=10^{-5}$). Training 1B model on 80B tokens, we compared these schedules across various LR and BS.
@@
Fig. 3 presents comparative heatmaps of the final training loss. We observe that compared to using a fixed final learning rate, setting it as $\mathrm{lr}_{\max}/10$ shows distinct optimal hyperparameter points and an overall left-skewed distribution of suboptimal learning rate and batch size combinations. We analyze that this is because, for the relatively high peak learning rates, conventional schedules result in disproportionately large minimum learning rates, which adversely affects the final stages of training and prevents the loss from converging to better local optima. As further illustrated in Fig. 1, aside from Porian Law, which converges the $\mathrm{lr}_{\min}$ to a sufficiently small value, the optimal learning rates calculated by other traditional learning rate decay schedules all exhibit varying degrees of a left-skew issue.
@@
This aligns with advanced training practices which suggest that the minimum learning rate significantly impacts the loss. This phenomenon is unfavorable for fitting our scaling laws, and in practice, it is generally preferred to keep the $\mathrm{lr}_{\min}$ fixed at a relatively low value. So we adopt the fixed final learning rate strategy in our subsequent experiments.
3.3.3 Evaluation Metric Consistency
Figure 4: Contour plots of training loss (left) and validation loss (right) . Both plots share the global minimum () at batch size 393,216 and learning rate of 0.001950.
As described in Chinchilla hoffmann_training_2022 , smooth training loss is considered an unbiased estimate of validation loss for simplicity. We operate under this same setting and supplement our investigation with experimental analysis. As shown in Fig. 4, for the case where the smooth training loss converges to the optimal value of $2.279$ ( as indicated by the solid red-framed line in Fig. 4 right) , the corresponding LR and BS are $1.95\times 10^{-3}$ and $393,216$ respectively. This is the same as the position of the LR and BS corresponding to the validation loss converging to the optimal value of $2.038$ (as indicated by the solid red-framed line in Fig. 4 left). Moreover, the overall trend of how the smooth training loss deviates from the optimal value with varying learning rates and batch sizes (as shown by the patterns in Fig. 4 right) closely mirrors the corresponding variations observed in validation loss measurements. This alignment demonstrates that the smooth training loss provides consistent optimization guidance for learning rate and batch size selection, matching the parameter configurations that would be obtained through direct validation loss evaluation.
3.4 Fitting Hyperparameter Scaling Laws